OpenStudy (anonymous):

**I give Medals** There are twenty coins in all pennies and nickels only the total value is 48 how many pennies are there and how many nickels??

OpenStudy (anonymous):

i tried so hard but in the end it dosent even matter life is like a art project you can make it as beautiful or ugly as you want make it good make it bad no one will care besides the person who loves you for who you are no matter your life style how you look all you need is true love and they wont care the way you look

OpenStudy (anonymous):

7 nickles and 13 pennies.

OpenStudy (anonymous):

7 nickels and 13 pennys

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

I think it's 7 nickles and 13 pennies

OpenStudy (anonymous):

noooo its 7 and 14

OpenStudy (anonymous):

7 Nickels 13 pennies

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@V_for_VENDETTA what are u talking about i just need help with the question not a life story

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

u copied my answer

OpenStudy (anonymous):

@kyrea lol

OpenStudy (anonymous):

music is right he needs the medal he said it first

OpenStudy (anonymous):

what = 48?

OpenStudy (anonymous):

actually i replied first sooooo

OpenStudy (anonymous):

ok so how did u get that answer @myronnapora

OpenStudy (anonymous):

well 7*5 is 35 then 35+13 is 48

OpenStudy (anonymous):

oka he is right yea yall are wrong its 48

TheSmartOne (thesmartone):

Please do not give direct asnwers. Take the time to read the CoC. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

who asked u to come here

TheSmartOne (thesmartone):

The way to solve your question. p is pennies n is nickels p+n=20 .01p+.05n=.48

OpenStudy (anonymous):

right thats what im saying @myronnapora

TheSmartOne (thesmartone):

@myronnapora The CoC states to be nice http://openstudy.com/code-of-conduct Open Study is an online thing and anybody cna come on the website and click on a question.

TheSmartOne (thesmartone):

And you can solve those equations by graphing, substituting, or eliminating.

OpenStudy (anonymous):

im being nice

OpenStudy (anonymous):

can someone just help god

OpenStudy (skullpatrol):

What is it that you are having trouble understanding?

OpenStudy (skullpatrol):

@kyrea ?

OpenStudy (skullpatrol):

There are twenty coins in all pennies and nickels only the total value is 48 how many pennies are there and how many nickels??

jabez177 (jabez177):

i think 9 nickels and 3 pennies

OpenStudy (anonymous):

how did they get that answer @skullpatrol

OpenStudy (skullpatrol):

What is the value of a penny?

OpenStudy (skullpatrol):

What is the value of a nickel?

OpenStudy (anonymous):

never mind dude dont need your help @LIGHT_EM_UP can u help

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hey what do you need help with

OpenStudy (anonymous):

Whtat does chu need help wit

OpenStudy (anonymous):

@kyrea what is the question

OpenStudy (skullpatrol):

There are twenty coins in all pennies and nickels only the total value is 48 how many pennies are there and how many nickels??

OpenStudy (anonymous):

9 and 3 pennies

OpenStudy (anonymous):

because 5x9 = 45 45 + 3 = 48

OpenStudy (anonymous):

thx so much ^^that helped alot @LIGHT_EM_UP

OpenStudy (anonymous):

9 nickles sowwy

OpenStudy (skullpatrol):

Let p = the number of pennies then 20 - p = the number of nickels, because p + n = 20.

OpenStudy (skullpatrol):

(5 cents/coin)*(number of nickels) + (1 cent/coin)*(number of pennies) = 48

OpenStudy (anonymous):

I did it the easy way dude stop trying to make it complicated cus its not

OpenStudy (skullpatrol):

5(20-p) + p = 48 100 - 5p + p = 48 52 = 4p p = 13 20 - 13 = 7 nickels

OpenStudy (skullpatrol):

13 pennies and 7 nickels Answer

OpenStudy (anonymous):

dude 9 nickles equals 45 3pennies to add to that makes it 48 retake the 6th grade dude

OpenStudy (skullpatrol):

@LIGHT_EM_UP dude,$$\Huge 9 + 3 = 12 \neq 20$$

OpenStudy (skullpatrol):

$$\Huge\text{There are twenty coins in all}$$ pennies and nickels only the total value is 48 how many pennies are there and how many nickels??